The demand function for product X is: Qd = 600 – 20Px + 0.02Y – 5Pr The supply function is: Qs = –300 + 10Px
Where:
Qd = the quantity of X demanded Qs = the quantity of X supplied Px = the price of product X Y = the average consumer income Pr = the price of the related product

If Y=35000 and Pr is R20, draw the demand curve
1
Expert's answer
2019-03-14T18:25:40-0400

Qd = 600 – 20Px + 0.02Y – 5Pr, Qs = –300 + 10Px 

If Y=35000 and Pr is R20, then the demand function is: Qd = 600 - 20Px + 700 - 100 = 1200 - 20Px, Px = 60 - 0.05Qd.

So, the demand curve is a downward-sloping line.